# Differentially Private Clustering in Data Streams

### v2

@@ Abstract
The streaming model is an abstraction of computing over massive data streams, which is a popular way of dealing with large-scale modern data analysis. In this model, there is a stream of data points, one after the other. A streaming algorithm is only allowed one pass over the data stream, and the goal is to perform some analysis during the stream while using as small space as possible.
Clustering problems (such as $k$-means and $k$-median) are fundamental unsupervised machine learning primitives, and streaming clustering algorithms have been extensively studied in the past. However, since data privacy becomes a central concern in many real-world applications, non-private clustering algorithms are not applicable in many scenarios.
In this work, we provide the first differentially private streaming algorithms for $k$-means and $k$-median clustering of $d$-dimensional Euclidean data points over a stream with length at most $T$ using $\operatorname*{poly}(k,d,\log(T))$ space to achieve a constant multiplicative error and a $\operatorname*{poly}(k,d,\log(T))$ additive error. In particular, we present a differentially private streaming clustering framework which only requires an offline DP coreset or clustering algorithm as a blackbox. By plugging in existing results from DP clustering [29, 42], we achieve (1) a $(1+\gamma)$-multiplicative approximation with $\tilde{O}_{\gamma}(\operatorname*{poly}(k,d,\log(T)))$ space for any $\gamma>0$, and the additive error is $\operatorname*{poly}(k,d,\log(T))$ or (2) an $O(1)$-multiplicative approximation with $\tilde{O}(k^{1.5}\cdot\operatorname*{poly}(d,\log(T)))$ space and $\operatorname*{poly}(k,d,\log(T))$ additive error.
In addition, our algorithmic framework is also differentially private under the continual release setting, i.e., the union of outputs of our algorithms at every timestamp is always differentially private.
@@ 1 Introduction
In real-world applications, a major challenge in dealing with large-scale data is that the entire datasets are too large to be stored in the computing system. The need to address this challenge and the success of large-scale systems (such as Spark Streaming [44]) that process data in streams have driven the study of the streaming model, which is introduced by a seminal work [3]. In this model, there is a stream of data points. During the stream, a data point arrives at each timestamp, and a streaming algorithm can only access these data points by a single pass. The goal is to output an (approximate) solution to a problem with respect to the set of data points that are arrived while using as small space as possible . If an algorithm is required to output at every timestamp when a new data point arrives, then it is called the continual release setting. Otherwise, the algorithm only needs to output at the end of the stream, which is called non-continual release or the one-shot setting. The streaming model has attracted a lot of attention from different areas in the past decades. In particular, streaming clustering algorithms have been extensively studied by the clustering literature.
Clustering is an essential primitive in unsupervised machine learning, and its geometric formulations, such as $k$-means and $k$-median, have been studied extensively, e.g., [4, 11, 32, 12, 13, 5, 40, 36, 2]. Euclidean $k$-clustering problem is stated as follows. Given a point set $P\subset\mathbb{R}^{d}$, the goal is to output a set of centers $C\subset\mathbb{R}^{d}$ with size $|C|=k$ such that the $k$-clustering cost $\textsf{cost}(C,P)=\sum_{p\in P}\min_{c\in C}\|p-c\|_{2}^{z}$ is minimized, where $z=1$ and $z=2$ correspond to $k$-median and $k$-means problem respectively. There is a long list of work (e.g., a subset includes [32, 31, 14, 26, 8, 18, 20]) studying $k$-means and $k$-median problem in the streaming setting. In the streaming $k$-clustering problem, there is a stream of points in $\mathbb{R}^{d}$, and the goal is to output a set of $k$ centers at each timestamp $t$, and minimize the $k$-clustering cost with respect to data points arrived before $t$. The state-of-the-art result is achieved by [20] which uses $\tilde{O}\left(\frac{kd}{\gamma^{2}}\right)\cdot\min\left(\frac{1}{\gamma^{z}},k\right)\cdot\operatorname*{poly}(\log\log T)$ space to obtain a $(1+\gamma)$-approximation with probability $0.9$ at the end of the stream. However, none of these algorithms are private, which means they are not applicable when the dataset involves personal information and privacy is considered to be a major concern in real-world applications.
@@
In this work, we present the first DP streaming algorithms for Euclidean $k$-means and $k$-median clustering using $\operatorname*{poly}(k,d,\log(T))$ space to achieve an $O(1)$-multiplicative error and a $\operatorname*{poly}(k,d,\log(T))$-additive error. In addition, our algorithms are DP under the continual release setting, which means that the union of all historical outputs of our algorithm is DP. Note that any DP algorithm under the continual release setting is always DP under the one-shot setting.
1.1 Computation Model and Differential Privacy
In this section, we formally define differential privacy and the streaming model under continual release setting. The input is a stream of points $x_{1},x_{2},\cdots,x_{T}\in\mathbb{R}^{d}$, where each $x_{i}\in\mathbb{R}^{d}$ satisfies $\|x_{i}\|_{2}\leq\Lambda$, i.e., we assume all input points are within a ball of radius $\Lambda$. Streams ${\mathcal{S}}=(x_{1},\ldots,x_{T})$ and ${\mathcal{S}}^{\prime}=(x^{\prime}_{1},\ldots,x^{\prime}_{T})$ are neighboring if there exists at most one timestamp $t^{*}\in[T]$ for which $x_{t^{*}}\neq x^{\prime}_{t^{*}}$ and $x_{t}=x^{\prime}_{t}$ for all $t\neq t^{*}$. In this paper, we consider every streaming algorithm ${\mathcal{A}}$ under the continual release setting, i.e., the entire output of ${\mathcal{A}}$ is $(s_{1},s_{2},\cdots,s_{T})$ where $s_{t}$ is the output of ${\mathcal{A}}$ at timestamp $t$ with respect to the data arrived no later than $t$. If we do not specify the timestamp, then the output of ${\mathcal{A}}$ indicates the entire output of ${\mathcal{A}}$ over the stream.
@@ 1.1 Computation Model and Differential Privacy
A randomized algorithm ${\mathcal{A}}$ is $(\varepsilon,\delta)$-DP if for every pair of neighboring datasets $D\sim D^{\prime}$, and for all sets ${\mathcal{S}}$ of possible outputs, we have that $\Pr[{\mathcal{A}}(D)\in{\mathcal{S}}]\leq e^{\varepsilon}\Pr[{\mathcal{A}}(D^{\prime})\in{\mathcal{S}}]+\delta$. When $\delta=0$ we simply say that the algorithm is $\varepsilon$-DP.
@@ 1.2 Our Results
We present a general framework for DP $k$-clustering in the streaming setting which utilizes an offline DP semicoreset (see Definition 3) algorithm as a black-box. Using existing results from the DP clustering literature, the cost of the resulting clustering output by our framework achieves (1) a $(1+\gamma)$-multiplicative error with space complexity having a $\operatorname*{poly}(k)$ dependency — using the DP coreset algorithm from [29] (see Theorem 1), or (2) an $O(1)$-multiplicative error with space complexity having a $k^{1.5}$ dependency — using the DP clustering algorithm from [42] (see Theorem 2). We use the term semicoreset as a generalization for any coreset that has an additive error proportional to the optimal cost. Notably any DP coreset is also a DP semicoreset, and [17] show how to obtain a DP semicoreset from an offline DP $k$-clustering algorithm which we will use as a blackbox in the sequel to achieve better space bounds in the case of Theorem 2.
@@
We state our results for $k$-means in the sequel, but note that our results easily generalize to $k$-median. In the following results, we assume we are given a non-DP algorithm in the offline setting that can compute a $(1+\gamma)$-approximation to $k$-means (e.g. [41]). As is standard in DP clustering literature, we assume $\Lambda$ is an upper bound on the diameter of the space of input points.
Theorem 1 (Informal version of Corollary 16).
@@
Given dimension $d$, clustering parameter $k$, $(1+\gamma)$-approx non-DP coreset algorithm $(\gamma\in(0,0.5))$ [18], $\varepsilon$-DP coreset algorithm from [29], and $(1+\gamma)$-approx non-DP $k$-means clustering algorithm [41]. Let ${\mathcal{S}}$ be the stream of input points in Euclidean space with length at most $T$. There exists a streaming algorithm ${\mathcal{A}}^{\prime}$ for $k$-means that outputs a set of $k$ centers ${\mathcal{C}}_{\hat{{\mathcal{Y}}}}$ at every timestep $t\in[T]$ such that
@@
    (Accuracy) With high probability,
    $\mathsf{cost}({\mathcal{C}}_{\hat{{\mathcal{Y}}}},{\mathcal{S}})$ $\leq(1+\gamma^{\prime})\mathsf{cost}({\mathcal{C}}^{opt}_{\mathcal{S}},{\mathcal{S}})+O_{\gamma}(1)\cdot V^{\prime}(d,k,\varepsilon,T,\Lambda)$ (1)
    where ${\mathcal{S}}$ denotes the set of all points given by the stream before timestamp $t$, ${\mathcal{C}}^{opt}_{\mathcal{S}}$ is the optimal solution, $\gamma^{\prime}=2\gamma+\gamma^{2}+\frac{2(1+\gamma)^{5}}{(1-\gamma)^{3}}=O(\gamma)$, and $V^{\prime}(d,k,\varepsilon,T,\Lambda)=\tilde{O}(\Lambda^{2}\cdot\operatorname*{poly}(k,d,\frac{1}{\varepsilon},\log(T)))$.
@@
    (Space) ${\mathcal{A}}^{\prime}$ consumes $\tilde{O}_{\gamma}\left(2^{O_{\gamma}(d^{\prime})}\cdot\operatorname*{poly}(k,d,\frac{1}{\varepsilon},\log(T))\right)$ space where $d^{\prime}=O_{\gamma}(\log k)$.
Theorem 2 (Informal version of Corollary 14).
Given dimension $d$, clustering parameter $k$, $(1+\gamma)$-approx non-DP coreset algorithm $(\gamma\in(0,0.5))$ [18], $(\varepsilon,\delta)$-DP k-means clustering algorithm from [42], and $(1+\gamma)$-approx non-DP $k$-means clustering algorithm [41]. Let ${\mathcal{S}}$ be a stream of input points in Euclidean space with length at most $T$. There exists a streaming algorithm ${\mathcal{A}}^{\prime}$ for $k$-means that outputs a set of centers ${\mathcal{C}}_{\hat{{\mathcal{Y}}}}$ at every timestep $t\in[T]$ such that

### v3

@@ Abstract
Clustering problems (such as $k$-means and $k$-median) are fundamental unsupervised machine learning primitives, and streaming clustering algorithms have been extensively studied in the past. However, since data privacy becomes a central concern in many real-world applications, non-private clustering algorithms may not be as applicable in many scenarios.
In this work, we provide the first differentially private algorithms for $k$-means and $k$-median clustering of $d$-dimensional Euclidean data points over a stream with length at most $T$ using space that is sublinear (in $T$) in the continual release setting where the algorithm is required to output a clustering at every timestep.
We achieve (1) an $O(1)$-multiplicative approximation with $\tilde{O}(k^{1.5}\cdot\operatorname*{poly}(d,\log(T)))$ space and $\operatorname*{poly}(k,d,\log(T))$ additive error, or (2) a $(1+\gamma)$-multiplicative approximation with $\tilde{O}_{\gamma}(\operatorname*{poly}(k,2^{O_{\gamma}(d)},\log(T)))$ space for any $\gamma>0$, and the additive error is $\operatorname*{poly}(k,2^{O_{\gamma}(d)},\log(T))$. Our main technical contribution is a differentially private clustering framework for data streams which only requires an offline DP coreset or clustering algorithm as a blackbox.
@@ 1 Introduction
In real-world applications, a major challenge in dealing with large-scale data is that the entire datasets are too large to be stored in the computing system. The need to address this challenge and the success of large-scale systems (such as Spark Streaming [49]) that process data in streams have driven the study of the streaming model, introduced by the seminal work [3]. In this model, there is a stream of data points — a data point arrives at each timestamp, and a streaming algorithm can only access these data points by a single pass. The goal is to output an (approximate) solution to a problem with respect to the set of data points that are arrived while using as small space as possible. If an algorithm is required to output at every timestamp when a new data point arrives, then it is called the continual release setting. Otherwise, the algorithm only needs to output at the end of the stream, which is called the one-shot setting. The streaming model has attracted a lot of attention from different areas in the past decades. In particular, streaming clustering algorithms have been extensively studied by the clustering literature.
Clustering is an essential primitive in unsupervised machine learning, and its geometric formulations, such as $k$-means and $k$-median, have been studied extensively, e.g., [4, 11, 35, 13, 14, 5, 45, 40, 2].
In the streaming $k$-clustering problem, there is a stream of points in $\mathbb{R}^{d}$, and the goal is to output a set of $k$ centers at each timestamp $t$, and minimize the $k$-clustering cost with respect to data points arrived before $t$. There is a long list of work (e.g., a subset includes [35, 34, 15, 29, 8, 19, 21]) studying $k$-means and $k$-median problem in the streaming setting. The state-of-the-art result is achieved by [21] which uses $\tilde{O}\left(\frac{kd}{\gamma^{2}}\right)\cdot\min\left(\frac{1}{\gamma^{z}},k\right)\cdot\operatorname*{poly}(\log\log T)$ space to obtain a $(1+\gamma)$-approximation with probability $0.9$ at the end of the stream. However, none of these algorithms are private, which means they are not applicable when the dataset involves personal information and privacy is considered to be a major concern in real-world applications.
@@
In this work, we present the first DP streaming algorithms for Euclidean $k$-means and $k$-median clustering using $\operatorname*{poly}(k,d,\log(T))$ space to achieve an $O(1)$-multiplicative error and a $\operatorname*{poly}(k,d,\log(T))$-additive error. Our algorithms are DP under the continual release setting. Note that any DP algorithm under the continual release setting is always DP under the one-shot setting.
1.1 Differential Privacy Model and Clustering Problem
We first formally define differential privacy and the streaming model under continual release setting and then define the clustering problem. The input is a stream of points $x_{1},x_{2},\cdots,x_{T}\in\mathbb{R}^{d}$, where each $x_{i}\in\mathbb{R}^{d}$ satisfies $\|x_{i}\|_{2}\leq\Lambda$, i.e., we assume all input points are within a ball of radius $\Lambda$. We work in the insertion-only setting, where points arrive sequentially but never leave the stream. That is, at time $t$ the dataset consists of all points $x_{1},\ldots,x_{t}$} seen so far, and future outputs must remain consistent with this growing prefix.
In this paper, we study streaming algorithms $\mathcal{A}$} under the continual release setting, i.e., the entire output of $\mathcal{A}$} is $(s_{1},s_{2},\cdots,s_{T})$ where $s_{t}$ is the output of $\mathcal{A}$} at timestamp $t$ with respect to the data arrived no later than $t$. We work under event-level privacy where the streams ${\mathcal{S}}=(x_{1},\ldots,x_{T})$ and ${\mathcal{S}}^{\prime}=(x^{\prime}_{1},\ldots,x^{\prime}_{T})$ are neighboring if there exists at most one timestamp $t^{*}\in[T]$ for which $x_{t^{*}}\neq x^{\prime}_{t^{*}}$ and $x_{t}=x^{\prime}_{t}$ for all $t\neq t^{*}$.
If we do not specify the timestamp, then the output of ${\mathcal{A}}({\mathcal{S}})$ indicates the entire output of $\mathcal{A}$} over the stream $\mathcal{S}$} at every timestamp.
@@ 1.1 Differential Privacy Model and Clustering Problem
A randomized algorithm $\mathcal{A}$} is $(\varepsilon,\delta)$-DP if for every pair of neighboring streams ${\mathcal{S}}\sim{\mathcal{S}}^{\prime}$, and for all sets $\mathcal{O}$} of possible outputs, we have that $\Pr[{\mathcal{A}}({\mathcal{S}})\in{\mathcal{O}}]\leq e^{\varepsilon}\Pr[{\mathcal{A}}({\mathcal{S}}^{\prime})\in{\mathcal{O}}]+\delta$. When $\delta=0$ we simply say that the algorithm is $\varepsilon$-DP.
For points $x,y\in{\mathbb{R}}^{d}$, we let $d(x,y)=\|x-y\|_{2}$ be the Euclidean distance between $x$ and $y$. Given a set $\mathcal{C}$}, we define $d(x,C):=\min_{c\in C}d(x,c)$.
For a set of centers $C$, we define the cost of clustering for the set $\mathcal{S}$} w.r.t. $C$ as
$\textsf{cost}(C,{\mathcal{S}})=\sum_{x\in{\mathcal{S}}}d^{z}(x,C)$
where $z=1$ for $k$-median, and $z=2$ for $k$-means.
Our goal in DP clustering is to produce a set of $k$ centers $C_{\mathcal{S}}$ for input stream $\mathcal{S}$} such that (1) $C_{\mathcal{S}}$ is $(\varepsilon,\delta)$-DP wrt $\mathcal{S}$}, and (2) $\textsf{cost}(C_{\mathcal{S}},{\mathcal{S}})\leq\alpha\cdot\textsf{cost}(C^{opt}_{\mathcal{S}},{\mathcal{S}})+\beta$.
@@ 1.2 Our Results
Before presenting our main results, we first define some useful terminology. The notation $O_{x}(\cdot)$ ignores factors involving $x$. We use the term semicoreset to describe a relaxed version of a coreset that allows an additive error proportional to the optimal cost. This notion is not intended as a canonical definition, but rather as a convenient abstraction that simplifies the presentation of our results. In particular, it enables a unified statement of Theorem 1 that can work using both static DP coreset algorithms and static DP clustering algorithms as a black-box.
Definition 2 ($(\kappa,\eta_{1},\eta_{2})$-semicoreset).
Given a point set $P$ in ${\mathbb{R}}^{d}$, a (multi)set $Q$ is a $(\kappa,\eta_{1},\eta_{2})$-semicoreset of $P$ for $k$-clustering ($k$-means or $k$-median) if
for any set of $k$-centers $C\subseteq{\mathbb{R}}^{d}$.
In particular, a coreset that has additive error that is not proportional to the optimal cost is simply a $(\kappa,0,\eta_{2})$-semicoreset. We also abuse notation and refer to a coreset with no additive error, i.e., $(\kappa,0,0)$-semicoreset as a $\kappa$-coreset.
@@
Given an offline DP $k$-clustering algorithm one can obtain an offline DP semicoreset by using a transformation (stated in Theorem 11) by [18] — we use this transformation as a blackbox in the sequel. We state our results for $k$-means, but note that our results easily generalize to $k$-median. As is standard in DP clustering literature (e.g.,[47]), we assume $\Lambda$ is an upper bound on the diameter of the space of input points. For ease of presentation, we assume $\Lambda=1$ in this part.
Our main result is a general framework for DP $k$-clustering in the streaming setting which utilizes an offline DP semicoreset algorithm as a black-box (see Theorem 1). Using existing results from the DP clustering literature, the cost of the resulting clustering output by our framework achieves (1) an $O(1)$-multiplicative error with space complexity having a $k^{1.5}$ dependency — using the DP clustering algorithm from [47] (see Theorem 2), or (2) a $(1+\gamma)$-multiplicative error with space complexity having a $\operatorname*{poly}(k)$ dependency — using the DP coreset algorithm from [32] (see Theorem 3).
We emphasize that our work establishes the theoretical feasibility of sublinear-space DP clustering in streams—a fundamental question that was previously open. As offline DP clustering algorithms achieve better additive error bounds, our black-box approach immediately inherits these improvements, potentially leading to more practical streaming variants.
We assume we are given a non-DP algorithm in the offline setting that can compute a $(1+\gamma)$-approximation to $k$-means (e.g. [46]) in the following statements.
Theorem 1 (Main).
@@
Given dimension $d$, clustering parameter $k$, arbitrary parameter $C_{M}$, a non-DP $(1+\gamma)$-coreset algorithm, an $(\varepsilon,\delta)$-DP $(\kappa,\eta_{1},\eta_{2})$-semicoreset algorithm $\mathcal{A}$} that outputs a semicoreset of size $SZ_{\mathcal{A}}(\cdot)$ and using space $S_{\mathcal{A}}(\cdot)$. Then, there exists a streaming algorithm $\mathcal{A}$$^{\prime}$ for $k$-means that outputs a set of centers $^{\mathcal{C}}_{\hat{{\mathcal{Y}}}}$ at every timestep $t\in[T]$ such that
  1. 1.
    (Privacy) $\mathcal{A}$$^{\prime}$ is $(3\varepsilon,\delta)$-DP under the continual release setting.
  2. 2.
    (Accuracy) With probability $1-\frac{1}{T^{2}}-\frac{1}{k^{O(\operatorname*{poly}(k)}}$,
    $\mathsf{cost}({\mathcal{C}}_{\hat{{\mathcal{Y}}}},{\mathcal{S}})$ $\leq{\kappa}\cdot\left((1+\gamma)C_{M}+\eta_{1}+\frac{(1+\gamma)^{4}}{(1-\gamma)^{3}}(\kappa+C_{M})\right)\cdot\mathsf{cost}({\mathcal{C}}^{opt}_{\mathcal{S}},{\mathcal{S}})+V(d,k,\varepsilon,\delta,T,\gamma)$
    where $\mathcal{S}$} denotes the set of all points given by the stream before timestamp $t$, ${\mathcal{C}}_{{\mathcal{S}}}^{opt}$ is the optimal solution, and $V(d,k,\varepsilon,\delta,T,\gamma)=\tilde{O}_{\gamma}(\kappa(\frac{C_{M}k^{2}}{d}+{kM})\frac{1}{\varepsilon}\operatorname*{poly}\log(T))$.
  3. 3.
    (Space) $\mathcal{A}$$^{\prime}$ consumes $\tilde{O}_{\gamma}(S_{\mathcal{A}}(M)+SZ_{\mathcal{A}}(M)+M+k\operatorname*{poly}\log(\frac{T}{M}))$ space.
where $M=O(\frac{d^{3}\eta_{2}}{C_{M}})$.
Remark.
The space complexity achieved by our framework depends on the additive error (i.e., the term $\eta_{2}$) of the offline DP $(\kappa,\eta_{1},\eta_{2})$-semicoreset algorithm used as a black-box. In particular, if there exists a DP semicoreset algorithm which achieves a better additive error in the offline setting, then our framework immediately gives a DP clustering algorithm with better space bounds in the continual release setting.
Remark.
The parameter $C_{M}$ in Theorem 1 is assigned an appropriate value depending on the error of the offline DP semicoreset algorithm used as a black-box in order to obtain the desired multiplicative approximation of the resulting algorithm $\mathcal{A}$$^{\prime}$. For e.g., we can set $C_{M}=O(1)$ to obtain an $O(1)$-multiplicative approximation in Theorem 2, and we can set $C_{M}=\gamma/2$ to obtain a $(1+O(\gamma))$-multiplicative approximation in Theorem 3.
We first present Theorem 2 which is the result of applying our framework to the DP $k$-means algorithm given by [47] in the offline setting that achieves an $O(1)$ multiplicative error. We note that [47] gave a DP coreset algorithm which has an additive error with $\operatorname*{poly}(k)$ dependency whereas their DP $k$-means algorithm achieves an additive error of $k^{1.5}$. This is why we choose to use their DP $k$-means algorithm and compute a DP semicoreset albeit with a slight loss in privacy to save on space (in terms of $k$) instead of directly using their DP coreset algorithm.
Theorem 2.
Given dimension $d$, clustering parameter $k$, privacy parameters $\varepsilon,\delta$, approximation parameter $\gamma>0$. There exists a streaming algorithm $\mathcal{A}$$^{\prime}$ for $k$-means that outputs a set of $k$ centers $^{\mathcal{C}}_{\hat{{\mathcal{Y}}}}$ at every timestep $t\in[T]$ such that
  1. 1.
  2. 2.
    (Accuracy) With probability $1-\frac{1}{T^{2}}-\frac{1}{k^{O(\operatorname*{poly}(k))}}$,
    $\mathsf{cost}({\mathcal{C}}_{\hat{{\mathcal{Y}}}},{\mathcal{S}})$ $\leq O_{\gamma}(1)\cdot\mathsf{cost}({\mathcal{C}}^{opt}_{\mathcal{S}},{\mathcal{S}})+V(d,k,\varepsilon,\delta,T)$
    where $\mathcal{S}$} denotes the set of all points given by the stream before timestamp $t$, ${\mathcal{C}}^{opt}_{\mathcal{S}}$ is the optimal solution, and $V(d,k,\varepsilon,\delta,T)=\tilde{O}_{\gamma}((k^{2.5}d^{3.51})\cdot\operatorname*{poly}(\log(T),\log(\frac{1}{\delta}),\frac{1}{\varepsilon}))$.
  3. 3.
    (Space) $\mathcal{A}$$^{\prime}$ consumes $\tilde{O}((k^{1.5}\cdot d^{4.51})\cdot\operatorname*{poly}(\log(T),\log(\frac{1}{\delta}),\frac{1}{\varepsilon}))$ space.
Next, we present the result of applying our framework to the DP coreset for $k$-means of [32] which achieves a multiplicative error of $1+\gamma$ and an additive error of $O_{\gamma}(\frac{k^{2}2^{{O_{\gamma}}(d)}}{\varepsilon}\operatorname*{poly}\log(T))$ in Theorem 3.
Theorem 3.
Given dimension $d$, clustering parameter $k$, privacy parameter $\varepsilon$, approximation parameter $\gamma>0$. There exists a streaming algorithm $\mathcal{A}$$^{\prime}$ for $k$-means that outputs a set of $k$ centers $^{\mathcal{C}}_{\hat{{\mathcal{Y}}}}$ at every timestep $t\in[T]$ such that
@@
    (Accuracy) With probability $1-\frac{1}{T^{2}}-\frac{1}{k^{O(\operatorname*{poly}(k))}}$,
    $\mathsf{cost}({\mathcal{C}}_{\hat{{\mathcal{Y}}}},{\mathcal{S}})$ $\leq(1+\gamma)\mathsf{cost}({\mathcal{C}}^{opt}_{\mathcal{S}},{\mathcal{S}})+V(d,k,\varepsilon,T)$
    where $\mathcal{S}$} denotes the set of all points given by the stream before timestamp $t$, ${\mathcal{C}}^{opt}_{\mathcal{S}}$ is the optimal solution, $V(d,k,\varepsilon,T)=\tilde{O}_{\gamma}(\frac{k^{3}d^{3}2^{O_{\gamma}(d)}}{\varepsilon}\cdot\operatorname*{poly}\log(T))$.
@@
    (Space) $\mathcal{A}$$^{\prime}$ consumes $\tilde{O}\left(\operatorname*{poly}\left(\frac{2^{O_{\gamma}(d)}}{\varepsilon}\log(T),k,d\right)\right)$ space.
Remark.
We note that the exponential dependency in $d$ in Theorem 3 is a direct consequence of the additive error of the DP coreset having the same dependency in [32].
1.3 Related Work
In the offline setting, private clustering was first studied by [33], and [28]. In particular, [33] used the exponential mechanism to produce a pure DP polynomial-time algorithm that achieves $(O(1),\tilde{O}(k^{2}\Lambda))$-approximation in discrete spaces. However, their algorithm is highly inefficient in Euclidean space (see [47] for a detailed exposition). [6, 47] focused on designing an efficient polynomial time algorithm for clustering that achieves a constant (multiplicative) factor approximation in high-dimensional Euclidean space by adopting the techniques of [33] while maintaining efficiency. [28] introduced the notion of private coresets and a recent line of work has adopted their techniques to give clustering algorithms with better approximation guarantees and efficiency [31, 44]. [47] gave a pure DP $k$-means algorithm with $O(1)$-multiplicative error, [12] achieved the same multiplicative error but improved the additive error to $\tilde{O}(k\sqrt{d}\Lambda^{2}/\varepsilon)$. [42] and [32] gave pure DP clustering algorithms that achieve optimal multiplicative error of $(1+\gamma)w^{*}$, where $w^{*}$ is the best approximation ratio for non-private $k$-means clustering.
Concurrent Works.
[39] concurrently released results on differentially private $k$-means clustering in the continual release setting. Their work is mostly incomparable with ours — their main focus is to optimize the approximation ratio under the more general setting of continual observation with the insertion and deletion of points, but their algorithm does not optimize for space and requires storing all the input points in the stream. More recently [38] showed that a twenty year old greedy algorithm for clustering can be made DP and gives DP clustering algorithms for both the static and continual release setting that achieves the same multiplicative error as [32] and additive error that has a $\log^{1.5}(T)$ dependency. This algorithm also does not optimize for space and inherently stores all the points in the stream.
In contrast, we work in the insertion-only streaming setting and our main focus is to optimize the approximation ratio using low space, i.e., $\operatorname*{poly}\log(T)$, in the continual release setting. Note that we achieve the same multiplicative error as [39] in Theorem 3.
1.4 Our Techniques
Our techniques apply to both $k$-means and $k$-median clustering, but for simplicity, we explain our techniques for the problem of $k$-means. We first outline the challenges to designing a DP $k$-means clustering algorithm in the continual release setting and then discuss the main ideas behind our DP clustering framework.
Naive Merge and Reduce approaches fail.
A standard streaming technique for clustering is the Merge-and-Reduce framework [35, 1, 30]. The idea is to maintain coresets $C_{1},\dots,C_{L}$ for $L=\log T$ levels (using a static coreset algorithm as a black-box) such that their union forms a coreset of all points seen so far. [19] showed that any point set has a $(1+\gamma)$-coreset with size $k\cdot\operatorname*{poly}(1/\gamma)$, thus each level has capacity threshold $M=k\cdot\operatorname*{poly}(\log T)$. When a new point $x$ arrives, it is inserted into $C_{1}$. For $i\in[L]$, if some $C_{i}$ exceeds size $M$, we compress it into a $(1+1/\operatorname*{poly}(\log T))$-approximate coreset $C^{\prime}$, reset $C_{i}\leftarrow\emptyset$, and insert $C^{\prime}$ into $C_{i+1}$. This process repeats until all levels respect the size bound. By induction, each $C_{i}$ is a $(1+1/\operatorname*{poly}(\log T))^{i}$-coreset of its underlying input points, and hence $C_{1}\cup C_{2}\cup\cdots\cup C_{L}$ is a $(1+1/\operatorname*{poly}(\log(T)))^{L}=(1+O(\log(T))/\operatorname*{poly}(\log(T)))$-coreset of the entire input stream.
We can adapt the classical merge-and-reduce method to the DP setting as follows. At level $1$, whenever the buffer $C_{1}$ reaches size $M$, we run a DP $k$-means coreset routine to compress it. The resulting summary is inserted into $C_{2}$, and higher levels merge summaries using a non-private coreset algorithm. By the post-processing property, this preserves differential privacy while avoiding repeated privacy loss at higher levels. Note that the size checks that trigger compression at each level can be implemented privately with the Sparse Vector Technique (SVT) [25, 36] (see Appendix C for details).
The key point is that any DP coreset must introduce some additive error $\eta=\Omega(\Lambda^{2})$ [33, 28]. If we apply this framework to the entire stream, the additive error introduced at the base still accumulates across levels: each merge adds the error from its children. After $\log T$ levels, this leads to a total additive error of $T^{\Omega(1)}\cdot\Lambda^{2}$, which is too large to yield meaningful guarantees.
Our Approach.
Our main technical innovation is to design methods that prevent the additive error from growing exponentially, as in the naive implementation of the DP merge-and-reduce framework described above.
Rather than applying a single instance of the DP merge-and-reduce framework to the entire input stream, we first partition the space $\mathbb{R}^{d}$ so that nearby input points are grouped together, and apply multiple instances of the DP merge-and-reduce framework to disjoint groups in parallel. To do this, we compute a set of candidate centers forming a bicriteria approximation [^1] Here, a bicriteria approximation means that the algorithm may output more than $k$ centers while still approximating the $k$-means cost. for $k$-means, and then assign input points to groups based on these centers. Within each group, we run a DP merge-and-reduce framework: the first level uses a DP clustering or DP coreset algorithm, while higher levels are computed using a non-DP coreset algorithm. This produces a DP semicoreset (see Definition 2). The semicoreset abstraction makes our framework flexible, since it can incorporate either a DP clustering algorithm or a DP coreset algorithm as the base primitive in the merge-and-reduce process.
For every timestamp $t\in[T]$, on a high-level, our algorithm (algorithm 1) does the following —